OpenStudy (anonymous):

simplify 3/4i so far i know that the denominator cant have i in it. So in this case you would use the discriminant -4i. And multiply that on the top and bottom but im confused on how to multiply

OpenStudy (anonymous):

\[\frac{3}{4i}=\frac{3}{4i} \times \frac{i}{i} =\frac{3i}{4i^2}=\frac{3i}{4(-1)}=-\frac{3i}{4}=0.75i\]

OpenStudy (anonymous):

why would it be regular i not -4i?

OpenStudy (anonymous):

You will get the same thing by going that way as well.

OpenStudy (anonymous):

oh i see, thanks